OpenStudy (anonymous):

In the high jump, Fran's kinetic energy is transformed into gravitational potential energy without the aid of the pole. With that minimum speed must Fran leave the ground in order to lift her center of mass 1.3m and cross the bar with a speed of 2.1 m/s. Halp

OpenStudy (jennychan12):

conservation of energy. 1/2v^2 +gh = 1/2v^2 +gh. 1/2v^2 +(9.8)(1/3) = 1/2(2.1)^2 +0 solve for v.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

4.59m/s?

OpenStudy (jennychan12):

wait. sorry i think i might have screwed up. cuz i accidentally put a 1/3 instead of a 1.3 and when i did it, i got a negative number, and you cant take a square root of a negative number. so i think the equation is 1/2v^2 + 0 = 1/2(2.1)^2 +(9.8*1.3) i got 5.47 m/s sorry, my bad.